MICROWAVE INVESTIGATION OF SULFURIC ACID MONOHYDRATE

\underline{D. L. FIACCO}, S. W. HUNT, and K. R. LEOPOLD, Department of Chemistry, University of Minnesota, Minneapolis, MN 55455.

We report the first microwave spectroscopic investigation of the 1:1 complex of H_{2}O-H_{2}SO_{4} and several of its deuterated and ^{18}O containing isotopomers. The complex is prepared in situ via reaction of water and SO_{3} using a co-injection source in which H_{2}O vapor is introduced into the early stages of an Ar + SO_{3} expansion. Both a- and c- type spectra for fourteen isotopomers have been measured and are consistent in all cases with that of a near-prolate rotor with appreciable dipole moment components along the a- and c- inertial axes. The spectra of the isotopomers containing H_{2}^{16}O are complicated by internal motion of the water unit affecting both the a-type K_{−1} = 0 and several of the c-type transitions. The possible internal motions giving rise to the observed spectral splittings as well as the ground state structural parameters of the H_{2}O-H_{2}SO_{4} complex will be discussed and compared to recent DFT calculations.^{a}^{,}^{b}

aS. Re, Y. Osamura, K. Morokuma, J. Phys. Chem. A 103, 3535 (1999).

bH. Arstila, K. Laasonen, A. Laaksonen, J. Chem. Phys. 108, 1031 (1998).